# When Knockoffs fail: diagnosing and fixing non-exchangeability of Knockoffs

### v1

@@ 3       Compute the residual $\widehat{\epsilon}_{j}=X_{j}-f_{j}(\mathbf{X_{-j}})$
We can now state the Knockoffs validity result, which is proven in Appendix A:
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ the original data and $\mathbf{\tilde{X}}\in\mathbb{R}^{n\times p}$ Knockoffs built using Algorithm 1 with prediction model $f$. Assume that $\forall j\in\llbracket p\rrbracket$, the prediction error of $f_{j}$ converges to 00, i.e. that $\forall j\in\llbracket p\rrbracket,\left\|\widehat{\boldsymbol{\epsilon_{j}}}-\boldsymbol{\epsilon_{j}}\right\|_{2}^{2}\xrightarrow[n\to\infty]{}0$. Then, Knockoffs and original variables are asymptotically exchangeable in the number of samples $n$:
$\forall S\subset\llbracket p\rrbracket,\mathit{d}^{*}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\xrightarrow[\enskip n\to\infty\enskip]{}0$
Remark 2.
When using a Lasso model for $f$, the prediction error converges to 00 under standard assumptions described in Theorem 11.2 of Hastie, Tibshirani and Wainwright, 2015. In turn, Theorem 1 holds when using a Lasso model.
4 When Knockoffs fail: diagnosing non-exchangeability
A direct consequence of non-exchangeability is that Knockoff statistics of null variables ${\{X_{j}\text{ for }j\in H_{0}\}}$ are no longer symmetrical – yet this property is key to achieving error control via valid Knockoffs as in Candès et al., 2018. Said otherwise, non-exchangeability can make Knockoff importance scores non comparable with real variable importance scores which leads to bias in Knockoff statistics. This is illustrated in Figure 3 using a simulated data setup described in Section 5.
On the left panel, we display the inverse Cumulative Distribution Function (CDF) of Knockoff statistics of null variables, where Knockoffs come from an oracle, with known covariance. On the central panel, we display the inverse CDF in the same setup but with a covariance estimate that relies on the Graphical Lasso. On the right panel, we display the inverse CDF of statistics obtained using non-parametric Knockoffs built from the data. Knockoff statistics of oracle Knockoffs are nearly perfectly symmetric, as seen on the left panel. Notice that Knockoff statistics of null variables using data-derived Gaussian Knockoffs are not symmetric: they are skewed towards positive values, signalling that real and Knockoff importance scores are not comparable. Non-parametric Knockoffs nearly recover exact symmetry of null Knockoffs statistics.
Figure 3: Inverse CDFs of Knockoff statistics of null variables for oracle and data-derived Knockoffs. On the left panel, we display Knockoff statistics of null variables in an oracle simulation setup where perfectly valid Knockoffs have been built – i.e., using the true covariance of the data, computed using a massive amount of samples. Notice that the inverse CDF is practically symmetrical, which ensures validity of the inference. On the central panel, we display Knockoff statistics of null variables in the same setup but with a covariance estimate that relies on the Graphical Lasso. Notice that null Knockoffs statistics are not symmetrical: they are skewed towards positive values, signalling that real and Knockoff importance scores are not comparable. On the right panel, we display Knockoff statistics of null variables using non-parametric Knockoffs. Notice that symmetry is recovered, which ensures reliable inference. We use $n=500$ samples and $p=500$ variables.
A sufficient condition for non-exchangeability. Clearly, Knockoffs exchangeability does not hold if $\mathbf{x}$ is non-Gaussian while $\mathbf{\tilde{x}}$ is Gaussian. One can simply take $S=\llbracket p\rrbracket$; then $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}=(\mathbf{\tilde{x}},\mathbf{x})\stackrel{{\scriptstyle d}}{{\neq}}(\mathbf{x},\mathbf{\tilde{x}})$. More broadly, any Knockoff generation procedure that relies on unfulfilled assumptions on variable distribution or dependence structure fails to replicate the distribution of $\mathbf{x}$. This leads to a natural sufficient condition to diagnose non-exchangeability: if there exists a classifier that is able to accurately distinguish samples from $\mathbf{x}$ versus samples from $\mathbf{\tilde{x}}$, then $\mathbf{\tilde{x}}\neq\mathbf{{x}}$ in distribution and exchangeability is violated. This idea is related to the C2ST (Classifier Two-Sample Testing) literature (Gretton et al., 2012; Lopez-Paz and Oquab, 2016).
Classifier Two-Sample Testing for Knockoffs. Formally, we wish to test the null hypothesis $H_{0}:\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{=}}\mathbf{{x}}$ given $n$ samples of each distribution. Following Section 3 of Lopez-Paz and Oquab, 2016, we proceed by constructing the dataset
Then, the $2n$ samples of $\mathcal{D}$ are shuffled at random and and split into disjoint training and testing subsets $\mathcal{D}_{\text{tr }}$ and $\mathcal{D}_{\text{te }}$, where $\mathcal{D}=\mathcal{D}_{\text{tr }}\cup\mathcal{D}_{\text{te }}$ and $n_{\text{te }}:=\left|\mathcal{D}_{\text{te }}\right|$. Note that in practice, this split is performed several times to mitigate randomness, as in classical cross-validation.
Definition 3 (C2ST statistic, Lopez-Paz and Oquab, 2016).
$\hat{t}=\frac{1}{n_{\text{te }}}\sum_{\left(z^{i},l^{i}\right)\in\mathcal{D}_{\text{te }}}1\left[g\left(z^{i}\right)=l^{i}\right]$
Intuitively, if $\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{=}}\mathbf{{x}}$, the test accuracy $\hat{t}$ should remain near 0.5, corresponding to chance-level. Conversely, if $\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{\neq}}\mathbf{{x}}$ and that the binary classifier is able to unveil distributional differences between the two samples, the test classification accuracy $\hat{t}$ should be greater than chance-level. The procedure is summarized in Figure 4.
Figure 4: Using a Classifier Two-Sample Test to diagnose exchangeability problems in Knockoffs. We input both input data and Knockoffs to a binary classifier with two different labels – the goal is to check whether a classifier can distinguish between real variables and Knockoff variables. If the classifier’s accuracy is substantially above chance level, exchangeability is violated since the joint distributions of input data and Knockoffs are not equal.
Exchangeability violation by improper sample pairing. Note that the C2ST diagnostic tool only tests that Knockoffs are indeed sampled from the same distribution as the original data. In practice, to have valid Knockoffs, a consistent pairing between original and Knockoffs samples is also needed. Said otherwise, knowing how to sample from the original data distribution is not enough to build valid Knockoffs. We illustrate this point via a simple experiment in the same simulation setup described in Section 5. The experiment consists in shuffling $50\%$ of the sample pairings of valid Knockoffs. The middle panel reports the FDP for each simulation run in both scenarios (shuffled and non-shuffled). The left and right sketches schematically illustrate shuffling sample pairings in $2D$. Note that the C2ST diagnostic is invariant to this shuffling.
Figure 5: Exchangeability violation by improper sample pairing. In the left boxplot, we build valid Knockoffs via the Gaussian algorithm and perform inference as in (Candès et al., 2018). In the right boxplot, we use the same Knockoffs but shuffle $50\%$ of the sample pairings before inference. We repeat this experiment $10$ times. Note that, coherently with theory, the $FDR$ is indeed controlled in the left boxplot. However, in the right boxplot, error control is lost due to improper sample pairings. The left and right sketches schematically illustrate shuffling sample pairings in $2D$.
As seen in Figure 5, FDR control is lost when shuffling $50\%$ of the sample pairings of valid Knockoffs. This shows that, beyond equality of the distributions, a proper sample pairing is needed to obtain valid Knockoffs. To test this in practice, one may compute the optimal assignment between real samples and Knockoff samples e.g. via the Hungarian algorithm (Kuhn, 1955). If the resulting assignment doesn’t match the original one, then Knockoffs cannot be exchangeable.
5 Experimental results
We aim at showing the practical consequences of exchangeability violations on error control and diagnosis via C2ST statistics.
Simulated data setup. To assess the consequences of varying degrees of non-exchangeability with Gaussian Knockoffs, we design a simulation setup in which we control the difficulty to produce valid Knockoffs. At each simulation run, we first generate i.i.d. Gaussian data with a tensor structure $\mathbf{X}\in\mathbb{R}^{n\times a\times b\times c}$ with $a\times b\times c=p$. The idea is to generate $n$ samples in a 3D feature space of dimensions $(a,b,c)$. Then, we apply an isotropic smoothing kernel of width $w$ across these three dimensions, mimicking a smooth three-dimensional structure, and flatten the data to obtain $\mathbf{X}\in\mathbb{R}^{n\times p}$.
Then, we draw the true support $\boldsymbol{\beta}^{*}\in\{0,1\}^{p}$. The number of non-null coefficients of $\boldsymbol{\beta}^{*}$ is controlled by the sparsity parameter $s_{p}$, i.e. $s_{p}=\lVert\boldsymbol{\beta}^{*}\rVert_{0}/p$. The target variable $\mathbf{y}$ is built using a linear model:
where $\sigma=\lVert\mathbf{X}\boldsymbol{\beta}^{*}\rVert_{2}/(\text{SNR}\lVert\boldsymbol{\epsilon}\rVert_{2})$ controls the amplitude of the noise, SNR being the signal-to-noise ratio. We choose the setting $n=500,p=500,s_{p}=0.1,\text{SNR}=2$. To obtain $p=500$, we use $(a,b,c)=(10,10,5)$. Note that using $w=0$ is equivalent to sampling i.i.d. Gaussians since no smoothing is applied in this case. We vary the kernel width $w$ in the interval $[0,1.25]$ to parametrize the difficulty of the problem. This parameter is closely related to the level of correlation observed in the data; as the kernel width increases, the data becomes increasingly correlated.
In all settings, we use the Graphical Lasso for covariance estimation (Friedman, Hastie and Tibshirani, 2008). For each of $N$ simulations, we compute the empirical FDP of the Vanilla Knockoff selection set $S$ and the C2ST statistic averaged across 5-fold cross-validation:
$\widehat{FDP}(S)=\frac{|S\cap\mathcal{H}_{0}|}{|S|},\quad\hat{t}=\frac{1}{n_{\text{te }}}\sum_{\left(z^{i},l^{i}\right)\in\mathcal{D}_{\text{te}}}1\left[g\left(z^{i}\right)=l^{i}\right].$
We first benchmark all existing methods for building Knockoffs for continuous data enumerated in Section 3 using $w=0.5$. This value yields local correlations levels that are comparable to those observed in fMRI data.
Figure 6: C2ST diagnostic metric for varying number of variables and samples. In the left panel, the number of variables $p$ grows and $n=p$ at each point. Note that, when $p>100$ the C2ST accuracy is clearly above chance level for all methods apart from Gaussian Knockoffs and Non-parametric Knockoffs, signaling non-exchangeability in this regime. In the right panel, the number of variables is kept constant with $p$ set to $500$. The number of samples varies from $100$ to $2000$. Apart from Non-parametric Knockoffs, all other methods fail to produce valid Knockoffs, even for large values of $n$.
The left panel of Figure 6 shows that all methods except Gaussian and non-parametric Knockoffs fail when $p$ grows larger (using $n=p$). Amongst the state-of-the-art methods, Gaussian Knockoffs are the best candidate as they maintain a C2ST accuracy of at most $0.6$. The proposed approach maintains chance-level accuracy in all settings. Note that the data at hand is Gaussian – therefore, the non-exchangeability of Gaussian Knockoffs is necessarily due to the problem of covariance estimation. In contrast, the proposed non-parametric approach circumvents this difficult estimation problem by relying on regressions.
The right panel of Figure 6 shows that increasing the number of samples $n$ for a fixed $p$ does not fix the problem, as all methods apart from Non-parametric Knockoffs fail to provide valid Knockoffs. Interestingly, C2ST accuracy rises as $n$ grows larger for all methods apart from Non-parametric Knockoffs. Intuitively, one might expect that as $n$ grows larger, producing valid Knockoffs becomes easier and therefore that the C2ST accuracy should decrease. A possible interpretation of this result is the presence of two competing effects: while learning features of the underlying joint distribution of the data is indeed easier when $n$ grows, the number of training samples accessible to the classifier also grows as it is equal to $2n$. Therefore, the discriminating power of the classifier also improves when $n$ grows larger. From now on, we discard the VAE and Deep approaches because of their poor empirical performance.
Figure 7: FDP and C2ST diagnostic metric for varying smoothing. We use the 3D smoothing kernel width to parametrize the correlation present in the data, which in turn tunes the difficulty to produce valid Knockoffs. For Gaussian Knockoffs, the FDR is controlled only in the easiest settings, i.e. $w\in[0,0.5]$. Note that using $w=0.5$ represents a common setting encountered on real data: it indeed yields local correlations levels that are comparable to those observed in fMRI data. For $w>0.5$ the achieved FDR is substantially above the target FDR, and it grows as the smoothing increases. Note that the C2ST accuracy is clearly above chance level for $w>0.5$, signaling non-exchangeability in this regime. Non-parametric Knockoffs – defined in Section 3.3 – preserve error control in all regimes, which is consistent with C2ST accuracy remaining near chance level.
The left panel of Figure 7 shows that, for Gaussian Knockoffs, the FDR is controlled only in the easiest settings, i.e. $w\in[0,0.5]$. For $w>0.5$ the achieved FDR is substantially above the target FDR, and it grows as the smoothing increases. The $C2ST$ accuracy is clearly above chance level for $w>0.5$. By contrast, non-parametric Knockoffs maintain FDR control in all settings, while the $C2ST$ accuracy remains near chance level.
Semi-simulated data setup. We now turn to evaluating the proposed approach and diagnostic tool on real data. Following Blain et al., 2023; Nguyen, Thirion and Arlot, 2022, we use semi-simulated data to evaluate the proposed method with observed $\mathbf{X}$. We use a simulated response $\mathbf{y}$ to be able to compute the FDP. We consider a first functional Magnetic Resonance Imaging (fMRI) dataset $(\mathbf{X}_{1},\mathbf{y}_{1})$ on which we perform inference using a Lasso estimator; this yields $\boldsymbol{\beta}_{1}^{*}\in\mathbb{R}^{p}$ that we will use as our ground truth. Then, we consider a separate fMRI dataset $(\mathbf{X}_{2},\mathbf{y}_{2})$ for data generation. The point of using a separate dataset is to avoid circularity between the ground truth definition and the inference procedure. Concretely, we discard the original response vector $\mathbf{y}_{2}$ for this dataset and build a simulated response $\mathbf{y}_{2}^{sim}$ using a linear model:
where we set $\sigma$ so that $SNR=4$. We consider $7$ binary classifications problems like "gambling" (rewards vs loss) taken from the HCP dataset. For each of these classification problems, the dataset consists in $778\times 2=1556$ samples and $1000$ features. These features are obtained by averaging fMRI signals within a Ward parcellation scheme, which is known to yield spatially homogeneous regions (Thirion et al., 2014).
As we use $7$ datasets of HCP, we obtain $42=7\times 6$ possible pairs. Since we consider $\boldsymbol{\beta}_{1}^{*}$ as the ground truth, the FDP can be computed. Figure 8 shows False Discovery Proportion levels for 42 semi-simulated fMRI datasets based on HCP data for 5 different Knockoff-based inference methods, using either Gaussian Knockoffs or non-parametric Knockoffs. All methods – see Blain et al., 2023 for a detailed comparison – exhibit problematic False Discoveries proportions using default Gaussian Knockoffs (left panel) and Graphical Lasso-based covariance learning. The expeted behavior is recovered using non-parametric Knockoffs (right panel). As one could expect, the C2ST metric obtains high discrimination power ($0.95$ prediction accuracy) for the Gaussian Knockoffs (signalling non-exchangeability), but is at chance (accuracy of $0.51$) for non-parametric Knockoffs. Regarding the pairing condition, the Hungarian algorithm signals that both Gaussian and Non-parametric Knockoffs are optimally paired with original variables.
Figure 8: Empirical FDP on semi-simulated data for 42 contrast pairs using Gaussian vs Non-parametric Knockoffs. We use 7 HCP contrasts C0: "Motor Hand", C1: "Motor Foot", C2: "Gambling", C3: "Relational", C4: "Emotion", C5: "Social", C6: "Working Memory". We consider all 42 possible train/test pairs: the train contrast is used to obtain a ground truth, while the test contrast is used to generate the response. Inference is performed using the 5 methods considered in the paper and the empirical FDP is reported. Notice that error control of all methods is violated using default Gaussian Knockoffs (left panel) and recovered using Non-parametric Knockoffs (right panel). The C2ST metric is coherent in both cases, with an average accuracy of $0.95$ for the Gaussian Knockoffs (signalling non-exchangeability) and $0.51$ for the non-parametric Knockoffs.
6 Discussion
Knockoffs are a powerful and efficient method for controlled variable selection. This inference procedure allows performing conditional variable selection in one round of inference, without considering each variable individually as in Conditional Randomization Tests approaches.
The statistical guarantees provided by this method rely on the ability to construct valid Knockoffs - namely, Knockoffs and original variables must be exchangeable so that false positives can be demonstrably controlled. However, constructing valid Knockoffs based only on the available observations is not trivial.
The difficulty of this problem is particularly salient in contexts where i) the data at hand has a strong dependence structure and ii) when the number of variables $p$ is large. Note that these are common characteristics of high-dimensional variable selection problems where Knockoffs are relevant, such as fMRI brain mapping or genomic data analyses. These characteristics make the estimation of the joint distribution of the data increasingly difficult, whether done through covariance estimation (as in the Gaussian algorithm) or via deep learning based techniques. Our first contribution in this context is to show that existing generation procedures are unable to produce valid Knockoffs in such settings.
This paper also provides insight into the consequences of exchangeability violations: experiments on both real and simulated data show the important impact of departures from exchangeability on the reliability of Knockoff-based methods. These issues can lead to instability in error control for all Knockoff-based methods. To address this challenge in real-world scenarios, we have introduced a diagnostic tool that provides practitioners with the means to identify exchangeability problems. This tool relies on two parts: a classifier two-sample test which aims at detecting distributional differences between original and Knockoff variables, and a procedure that controls the proper pairing between original observations and their knockoff. Note that having a perfect generative model is not enough to obtain valid Knockoffs, as this does not yield samples paired with the original data. We therefore have two necessary (but not sufficient) conditions, for exchangeability violation.
As part of our efforts to mitigate the exchangeability problems associated with non-Gaussian data or poor covariance estimation, we propose an efficient alternative approach for constructing non-parametric Knockoffs. We prove theoretically that this approach produces valid Knockoffs, provided that the number of samples is large enough. Experiments on simulated data indicate that the Knockoffs we obtain are valid. However, the proposed approach has some limitations. In particular, a sufficient condition for exchangeability is still lacking, so that one may indeed encounter situations where supposedly valid Knockoffs lead to biased FDR guarantees. Another limitation in the non-parametric Knockoff estimation procedure is that the hypothesis that the learner is Bayes optimal leaves open the question of which learner to choose; this choice is obviously important in practice, both for the validity of the Knockoff inference and for the power of the method.
The code for the proposed diagnostic tool and alternative non-parametric Knockoffs algorithm is available at https://github.com/alexblnn/KnockoffsDiagnostics. Note that this algorithm is also less costly in terms of computing time than the Gaussian algorithm. Using a single CPU, a 3×3$\times$3 × speedup is achieved compared to the original algorithm. When leveraging 40 parallel CPUs, a 15×15$\times$15 × speedup is achieved on fMRI data and other large problems. Details about this benchmark are available in Appendix B. Since the proposed algorithm is by design robust to deviations from Gaussianity, it should be preferred in all situations.
Appendix A Proof of Non-parametric Knockoffs validity
Proof of Theorem 1. Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ the original data and $\mathbf{\tilde{X}}\in\mathbb{R}^{n\times p}$ Non-parametric Knockoffs built using Algorithm 1 with a given permutation of $\llbracket p\rrbracket$ denoted $\tau$. Denote $S\subset\llbracket p\rrbracket$ and $j\in\llbracket 2p\rrbracket$. We have, with $(f_{j},\boldsymbol{\widehat{\epsilon}}_{j})$ such that $X_{j}=f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}$:
$(\mathbf{X},\mathbf{\tilde{X}})_{j}=\left\{\begin{array}[]{l}f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}\text{ if }\quad j\in\llbracket p\rrbracket\\ f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{\tau{(j})}\text{ if }\quad j\in\llbracket p+1,2p\rrbracket\end{array}\right.$
and
$(\mathbf{X},\mathbf{\tilde{X}})_{swap(S),j}=\left\{\begin{array}[]{l}f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{\tau{(j})}\text{ if }\quad j\in S\cap\llbracket p\rrbracket\\ f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}\text{ if }\quad j\in S\cap\llbracket p+1,2p\rrbracket\\ (\mathbf{X},\mathbf{\tilde{X}})_{j}\quad\text{ else. }\end{array}\right.$
Let $\mu$ and $\nu$ two distributions of $\mathbb{R}^{2p}$. The usual 2-Wasserstein distance between $\mu$ and $\nu$ is defined as follows, with $\Theta(\mu,\nu)$ the set of distributions of $\mathbb{R}^{2p}\times\mathbb{R}^{2p}$ with marginals $\mu$ and $\nu$:
@@ Appendix A Proof of Non-parametric Knockoffs validity
$W_{2}(\mu,\nu):=\left(\inf_{\theta\in\Theta(\mu,\nu)}\int_{\mathbb{R}^{2p}\times\mathbb{R}^{2p}}||x-y||^{2}\mathrm{d}\theta(x,y)\right)^{1/2}$
It follows that $W_{2}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\leq W_{2}(\widehat{\boldsymbol{\epsilon}},\widehat{\boldsymbol{\epsilon}}_{\tau})\leq W_{2}(\widehat{\boldsymbol{\epsilon}},\boldsymbol{\epsilon})+W_{2}(\widehat{\boldsymbol{\epsilon}}_{\tau},\boldsymbol{\epsilon}_{\tau})$ with $\mathbf{\epsilon}$ the true noise vector such that:
$\forall j\in\llbracket p\rrbracket,\quad X_{j}=f^{*}_{j}(\mathbf{X_{-j}})+\boldsymbol{\epsilon}_{j}.$
[…]
$W_{2}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\xrightarrow[\enskip n\to\infty\enskip]{}0$
Since $W_{2}$ is a distance over distributions in $\mathbb{R}^{2p}$, this proves that Non-parametric Knockoffs are exchangeable asymptotically in the number of samples $n$.
Appendix B Computation time of non-parametric Knockoffs

### v2

@@ Appendix A Proof of Proposition 1
$W_{2}(P_{1},\widehat{P}_{1})$ $=\left(\underset{P_{\theta}\in\Theta(P_{1},\widehat{P}_{1})}{\mathrm{inf}}\int_{\mathbb{R}^{2p}\times\mathbb{R}^{2p}}(x-y)^{2}P_{\theta}(\mathrm{dx,dy})\right)^{\frac{1}{2}}$
$\leq\left(\int_{\mathbb{R}^{2p}}\bigg{(}\mathbb{E}\left[X_{1}\middle|X_{-1}\right]+\left(X^{\prime}_{1}-\mathbb{E}\left[X^{\prime}_{1}\middle|X^{\prime}_{-1}\right]\right)\right.$
$\quad\left.-f_{1}\left(X_{-1}\right)-\left(X^{\prime}_{1}-f_{1}\left(X^{\prime}_{-1}\right)\right)\bigg{)}^{2}P_{X}(\mathrm{dx})P_{X^{\prime}}(\mathrm{dx}^{\prime})\right)^{\frac{1}{2}}$
$=\left(\int_{\mathbb{R}^{2(p-1)}}\bigg{(}f^{*}_{1}\left(X_{-1}\right)-f^{*}_{1}\left(X^{\prime}_{-1}\right)\right.$
$\quad\left.-f_{1}\left(X_{-1}\right)+f_{1}\left(X^{\prime}_{-1}\right)\bigg{)}^{2}P_{X_{-1}}(\mathrm{dx}_{-1})P_{X^{\prime}_{-1}}(\mathrm{dx}^{\prime}_{-1})\right)^{\frac{1}{2}}$
$\leq\left(\mathbb{E}\left[\left(f^{*}_{1}\left(\mathbf{x}_{-1}\right)-f_{-1}\left(\mathbf{x}_{-1}\right)\right)^{2}\right]\right)^{\frac{1}{2}}$
$\quad+\left(\mathbb{E}\left[\left(f^{*}_{1}\left(X^{\prime}_{-1}\right)-f_{1}\left(X^{\prime}_{-1}\right)\right)^{2}\right]\right)^{\frac{1}{2}}.$
To conclude, we note that there is a need for uniform control. This is achieved by the limited complexity of the class of regressors (Donsker class) and its consistency. Under these assumptions, we apply Lemma 19.24 from Van der Vaart,, 2000. Therefore, there is uniform convergence, and the last two terms vanish.
For the $j$-th coordinate, we proceed similarly as before. By the induction hypothesis we have that $\tilde{X}_{l}|X_{-l},\tilde{X}_{1:l-1}\sim\mathcal{L}(X_{l}|X_{-l},\tilde{X}_{1:l-1})$ for each $l<j$. Therefore, given $(X^{\prime},\tilde{X}^{\prime}_{1:j-1})\overset{\mathrm{i.i.d}}{\sim}(X,\tilde{X}_{1:j-1})$, let $P_{j}$ be the distribution of $\mathbb{E}\left[X_{j}\middle|X_{-j},\tilde{X}_{1:j-1}\right]+\left(X^{\prime}_{j}-\mathbb{E}\left[X^{\prime}_{j}|X^{\prime}_{-j},\tilde{X}^{\prime}_{1:j-1}\right]\right)$ conditionally on $(X_{-j},\tilde{X}_{1:j-1})$. Analogously, let’s denote its empirical counterpart as $\widehat{P}_{j}:=f_{j}\left(X_{-j},\tilde{X}_{1:j-1}\right)+X^{\prime}_{j}-f_{j}\left(X^{\prime}_{-j},\tilde{X}^{\prime}_{1:j-1}\right)$. By Lemma 1, we have that $P_{j}$ is indeed the desired distribution, i.e. $\mathcal{L}(X_{j}|X_{-j},\tilde{X}_{1:j-1})$. Similarly as before, using the consistency of $f_{j}$ we conclude that $W_{2}(P_{j},\widehat{P}_{j})\xrightarrow[\enskip n\to\infty\enskip]{}0.$
Appendix B Nonparametric knockoffs with complex learners
Let’s denote $\mathbf{X}^{i}$ as the $i$-th row of the matrix $\mathbf{X}$, which corresponds to the $i$-th individual.
Input : $f,K$
1 Generate $B_{n}\in[1,K]^{n}$ by sampling uniformly from $[1,K]$ with replacement, and for $k\in[1,K]$, denote by $D_{k}:=\{i\mid B_{n,i}=k\}$ the set of indices in the $k$-th fold for $k\in[1,K]$ do
2       for $j\in[1,p]$ do
3            Fit a prediction model $f_{j,k}$ on $\{((\mathbf{X}^{i}_{-j},\bar{\mathbf{X}}^{i}_{1:j-1}),\mathbf{X}_{j}^{i})\}_{i\in\cup_{l\neq k}D_{l}}$
4             Compute the residual $\widehat{\boldsymbol{\epsilon}}_{j,k}=\{\mathbf{X}^{i}_{j}-f_{j,k}((\mathbf{X}^{i}_{-j},\tilde{\mathbf{X}}^{i}_{1:j-1}))\}_{i\in D_{k}}$
5             Sample $\text{Sample }\bar{\mathbf{X}}^{i}_{j}=f_{j,k}((\mathbf{X}^{i}_{-j},\bar{\mathbf{X}}^{i}_{1:j-1}))+\varepsilon_{i,j,k}$ with $\varepsilon_{i,j,k}$ sampled uniformly in $\widehat{\boldsymbol{\epsilon}}_{j,k}$ for $i\in\cup_{l\neq k}D_{l}$
6             Sample $\text{Sample }\tilde{\mathbf{X}}^{i}_{j}=f_{j,k}((\mathbf{X}^{i}_{-j},\tilde{\mathbf{X}}^{i}_{1:j-1}))+\varepsilon_{i,j,k}$ with $\varepsilon_{i,j,k}$ sampled uniformly in $\widehat{\boldsymbol{\epsilon}}_{j,k}$ for $i\in D_{k}$
7       end for
8
9 end for
10
Return $\mathbf{\tilde{X}}_{1:p}$
Algorithm 4 Cross-fitted sequential generation of nonparametric knockoffs by learning to predict $\mathbf{X}_{j}$ from $(\mathbf{X}_{-j},\tilde{\mathbf{X}}_{1:j-1})$ using a model $f_{j}$.
The proof of the validity of Algorithm 4 follows exactly the same reasoning as the one given in Appendix A, but it directly uses the consistency of the estimator to conclude, without requiring the uniform control needed to account for the dependency of the residuals on the trained regressors.
Appendix C Computation time of nonparametric knockoffs